# Intergroup violence in bursts or fizzles

### v9

@@ 3 Data and Methods
In studying violence, lab experiments lack the turmoil and emotional intensity of real violent confrontations due to their obligation to meet ethical standards. Field studies, in contrast, cannot be based on a random sample of participants or groups, yet they are invaluable for realistic view on violence [4]. We obtained 42 videos from websites such as YouTube, LiveLeak, and WorldStarHipHop using search terms with the English keywords “brawl,” “street fight,” and “assault.” This sample is random with respect to temporal unfolding and (sub)group size. Of these clips, 36 are from English-speaking countries (mainly the US and the UK, with one from Canada and one from India); five of the remaining clips are from the Netherlands, and one is from Colombia. We did not observe differences in relevant behavior related to the location of the recording. To keep distracting factors away from our analysis, we excluded clips with professional fighters, long range weapons, protective clothing, a referee, ambush attacks, or youths in a school yard. Most of our videos are recorded on phones by bystanders and are left-truncated; in all likelihood, there would have already been some turmoil that motivated bystanders to start filming. The shortest lasted 30 sec. and the longest was nearly 5 minutes (mean 101 sec.; s.d. 59 sec.). Out of a potential 2 x 42 groups, where the opponents in one analysis become the focal group in the next, 25 groups attacked a single individual rather than a group. Because a lone individual is unable to act collectively, this leaves 59 groups to examine; see the Supplementary Material for all case studies and elaborate examples of interpreting and coding. One group had 14 members, but all other groups were small, $102\leq n<$ 10 (mean 3.6). The smaller groups were simulated as cliques wherein everyone could see one another unless there were obstacles or de-escalators obstructing visual contact.
@@
We coded violence when force was used against another’s body (punching, slapping, kicking, hitting, stomping) and/or when another person’s body was forcefully moved. People may defect (not use violence) for all kinds of reasons and due to various causes. Yet, unconditional defectors have in common that they are (temporarily) inactive while others contribute. Among them we include de-escalators, as well as group members who were prevented to fight by de-escalators.
We subsumed the following behaviors of members of the opponent group under turmoil for the focal group: aggressing (including fighting gestures); pulling off clothing (jackets or vests); pulling up pants; pointing toward opponents; provocative gesturing with fingers or hands (as an invitation to engage); bending forward toward an opponent; approaching the focal group; encroaching (invading opponents’ personal space through using or damaging objects belonging to them); teasing, such as lightly hitting or ridiculing; and violence. Because stumbling and falling because vulnerability of opponents tends to agitate focal group members [15, 40, 62], we included these mishaps in our measure. It is likely that over relatively short time intervals, turmoil accumulates in participants’ minds. Therefore we calculated the level of turmoil from the beginning of the video until a focal group’s (first) maximum participation in violence. We did this by multiplying the duration of each instance of turmoil by the number of individuals involved, and adding up all these weighted instances. This scale should be improved in future studies, but our main prediction is independent of it.
@@ 4 Results
We confirm the finding of an earlier study [15] that fighting tends to start in small groups or in small subgroups of larger groups. As predicted, we found that small (sub)groups burst into action at lower levels of turmoil than larger groups [9], and lower levels are of course reached earlier. In our data, bursts developed in 13 (37%) of the 35 smallest groups (dyads and triads) and in 10 (42%) of the 24 larger groups. In bursts, the correlation between group size and turmoil is $0.53$. The size-turmoil relationship is slightly disturbed by dyads more often facing a larger opponent group and being therefore less likely to fight collectively than triads, which in turn are more robust in general [52]. In fact, larger groups always defeated smaller opponent groups (made them flee or worked them to the ground) with only one exception.
@@
The proportions of unconditional defectors in groups with bursts (mean = 0.19; s.d. = 0.21) and groups without bursts (mean = 0.49; s.d. = 0.27) are box-plotted in Fig. 5 (Welch test $t=4.796;P=6.411$ x $10^{-6}$; df = 54.76). The predicted critical threshold (vertical line in the figure) neatly separates the two boxplots, but one might question how robust this result is, given some noise in coding. There is a risk of missing a violent act (it can be very fast) or mislabeling a fast movement as violence, thus mixing up fighters and defectors, and a risk of misassigning individuals to groups. Even if we suppose there is a 10% chance of each of these errors (which corresponds to Krippendorff’s $\alpha=0.75$; see inter-coder reliability in the Supplementary Material), the result is hardly weakened (Welch test $t=3.879;P=0.0025$, averaged over 1000 simulation runs), hence the distinction between bursts and fizzles appears to be robust.
@@
Figure 5: The proportion of unconditional defectors and the predicted critical level (vertical line), which for our data separates the boxplots of bursts (1) and fizzles (0).
@@
The empirical distinction between bursts and fizzles is not perfect, and Fig. 5 shows that 15 groups out of 59 (25%) are categorized incorrectly. In 11 of these 15 groups, the time between the first violent act and last participant joining in was beyond our two-second interval, hence we did not count these collective actions as bursts. In simulations, lags between first instances of violence and the moment when at least half of the group is participating frequently occur when turmoil lingers at its critical level (Fig. 6), versus steadily increasing turmoil that yields a clean burst (Fig. 3). Lingering turmoil seems to have occurred in 4 cases, but we would need a better scale of turmoil than the cumulative scale we currently use to be sure. Far more important are local circumstances beyond the model. When re-examining the time lags beyond 2 seconds, we noticed various causes of delay: (1) group members were in a car and it took time to get out; (2) a focal group member stood at a distance from the fight on a slippery floor next to a pool; (3) a focal group member was beaten and seemed intoxicated by alcohol, hence responded slowly; (4) a focal group member was first pushed over and it took him time to get on his feet again; (5) opponents moved around so fast that it took time to hit one; (6) an opponent fell into thick bushes such that only one focal group member at a time could get to him; (7) group members were sitting in a train and needed time to get on their feet; (8) group members were constrained in a small room where they searched for a space to lash out; and, (9) group members were hindered by the mess of two opponents fighting on a collapsed table full of food. One can hardly blame the model for not incorporating these local circumstances that caused delays, and perhaps our 2 seconds limit should be applied with more care. Finally, two delays were caused by ambiguous behavior: (10) someone tried to de-escalate, was hit, and then attacked, and (11) someone tried to pull back his fighting friend (de-escalation) but attacked simultaneously.
@@
Figure 6: Proportion of cooperators ($N_{C}/n$) over time (or precisely, over Monte Carlo steps) in a fully connected network ($n=5$) with one unconditional defector, at $T=0.11\approx T_{c}$. Note an extended period of multiple short spikes of violence before all conditional cooperators join in. Similar lags occur without unconditional defectors.
@@
In the 4 out of 15 remaining groups, non-bursts were predicted whereas bursts occurred. Although for parsimony, we applied the same $S_{0}$ (and $\Delta$) for all conditional cooperators, some combatants ($i$) may actually have a higher value than the average, $S_{0,i}>\overline{S_{0}}$, which could explain these cases. In lab experiments, it is unlikely that the patterns of bursts and fizzles would have been discovered, but now they have been pointed out in the field, ethically acceptable intergroup conflicts in the lab may benefit from controlled circumstances, such as equal monetary payoffs for all cooperators, and may lead to refinements of the model.
@@ 4.1 Alternative explanation
Although in 21 out of 23 bursts, some degree of synchronization (10.8 on average) preceded collective violence, there were 18 cases in which synchronization (9.6 on average) was not followed by collective violence. In some of the latter cases, synchrony turned out to be a deceptive performance composed of blustering and aggrandizing without commitment to fighting. This does not imply that synchronization is unimportant (it probably is for solidarity [15]), but it does not predict collective violence.
5 Discussion
@@
The simple Ising model is a century old [31] and has been applied to a wide range of problems [44, 55], to which we add the dilemma of collective action. It explains cooperation parsimoniously, based on agitating stimuli without recourse to strong rationality, initiative takers, reputations, norms, feedback through selective incentives, or reliable information passing through the network. The Ising model elucidates the temporal unfolding of violence by predicting a critical threshold of unconditional defectors that distinguishes a burst of collective action from a fizzle, which is largely supported by the data. The model also explains why violent groups are often small or are small subgroups of larger groups despite greater risk. Small (sub)groups have a lower critical threshold of turmoil, and in a confrontation with opponents, lower levels are reached earlier. We also investigated whether synchronous action precedes violence, but we found that synchronization precedes both collective and solitary violence, and cannot predict either of these outcomes. However, synchronization may still be important to increase solidarity [18].
@@
This study has several limitations. Because we did not select any videos without violence, we cannot be certain that violence is caused by turmoil. When investigating the videos, however, we observed time and again that people reacted violently to turmoil, hence it still seems quite plausible that turmoil triggers violence. Moreover, the model predicts that when turmoil increases, small groups start cooperating earlier than large groups, which we were unable to falsify. Our measurements underestimated turmoil because the videos are left-truncated, and our data depend on camera angle and vision width, inevitably excluding instances of turmoil. Furthermore, there might be some error in our coding. If the truth were perfectly known and the coding corrected accordingly, the two boxplots in Fig. 5 might shift to a small degree, but this would would not affect our main result. In the future, a better scale of agitation should be developed. The scale should differentiate between domains (e.g., street violence versus grievances about politics), thus there should be multiple scales, and these should differentiate between effect sizes of different stimuli. For example, the impact of family members being killed will be much larger and last longer than that of a push, which might be quickly forgotten. Some protest movements are known to produce agitation by themselves, which is sometimes indicated by accelerated posting of online messages. Because in these cases, the inter-posting time intervals follow a certain pattern (Moore’s law), $T_{c}$ can be accurately predicted, well ahead of large street demonstrations [28]. Another limitation is in the Ising model itself. Despite predicting the threshold of unconditional defectors well, some of the groups ended up in the non-burst category due to local circumstances (ignored by the model) that caused delays beyond our two second limit. Furthermore, the Ising model does not predict the severity of violence. For future studies, it is important to expand the number and diversity of cases, and to develop better software than is currently available to automatically code videos. Finally, when individuals find themselves more often in similar situations, they will learn, which is easier in smaller groups where they have a larger influence on their payoff [11]. Some will change their strategy, and turn into unconditional defectors [2] who try to exploit other group members and maximize their individual payoff instead of maximizing the group’s payoff. For the model, this would require individuals updating their decision rules at subsequent Monte Carlo steps.
@@
Here we applied the Ising model to violence, but it can also be applied to other uncertain situations. Its dynamics are entirely consistent with temporal patterns of protests [24], which break out more often if (rumors say that) a government or its police are weakened [53, 58], analogous to vulnerable individuals in street fights. The model also seems applicable to spontaneous lynchings [6], and bystanders collectively helping victims under uncertainty [47]. In this first empirical application, we showed that it can explain the unfolding of street violence, and in all likelihood, more discoveries lay ahead. Extensions to norms (as external fields) and noise in actors’ information about others’ behavior have been explored in simulations [8].
@@
Taking a broader perspective, random noise can solve coordination problems, such as collective responses to opponents’ threats, by shaking a group loose from its suboptimal (e.g., non-cooperative) state [51], but it can also disturb an optimal state. In the Ising model, both can happen, depending on the amount of turmoil: when in a group, turmoil approaches $T_{c}$ from below, random noise in decision making facilitates cooperation, but at high levels, too much noise entails confusion.
Ethics
Data availability statement
All coded video data, the R script used to produce plots (Supplementary Material) from the data, and a Fortran script for simulations of the Ising model are available at https://osf.io/f25nq/
@@ 6 Data and compilation of data set
One concern about phone-recorded clips is that recorders start filming when the antagonism is already ongoing; our data are thus left-truncated to an unknown degree. We discarded footage that immediately started with physical violence. Another concern is that uploaded videos might be biased toward more spectacular cases. However, our sample shows variety in the forms and severity of violence, ranging from groups who engage in frenzied collective violence to incidents that involve just a few slaps. A third concern is whether the recording influenced the behavior of the assailants. Several videos showed more than one person recording the incident. Earlier work [30] noted that a host of research indicates that people do not change their behavior substantially in the presence of a camera and mainly focus on what they are doing. This is probably even more the case in antagonistic situations, and because young people are used to being filmed by phones. Additionally, when people do pay attention to the camera, which happened in just one of our clips, it is visible in the recording and can be incorporated in the analysis. (In our case, this concerned behavior coded as turmoil.)

### v10

@@ 4 Data and Methods
In studying violence, lab experiments lack the turmoil and emotional intensity of violent confrontations due to their obligation to meet ethical standards. Field studies, in contrast, cannot be based on a random sample of participants or groups, yet they are invaluable for realistic view on violence [4]. We obtained 42 videos from websites such as YouTube, LiveLeak, and WorldStarHipHop using search terms with the English keywords “brawl,” “street fight,” and “assault.” This sample is not random with respect to violence, but it is random with respect to temporal unfolding and (sub)group size. Of these clips, 36 are from English-speaking countries (mainly the US and the UK, with one from Canada and one from India); five of the remaining clips are from the Netherlands, and one is from Colombia. We did not observe differences in relevant behavior related to the location of the recording. To keep distracting factors away from our analysis, we excluded clips with professional fighters, long range weapons, protective clothing, a referee, ambush attacks, or youths in a school yard. Most of our videos are recorded on phones by bystanders and are left-truncated; in all likelihood, there would have already been some turmoil that motivated bystanders to start filming. The shortest lasted 30 sec. and the longest was nearly 5 minutes (mean 101 sec.; s.d. 59 sec.). Out of a potential 2 x 42 groups, where the opponents in one analysis become the focal group in the next, 25 groups attacked a single individual rather than a group. Because a lone individual is unable to act collectively, this leaves 59 groups to examine; see the Supplementary information for all case studies and elaborate examples of interpreting and coding. One group had 14 members, but all other groups were small, $102\leq n<$ 10 (mean 3.6). They were simulated as cliques wherein everyone could sense one another unless there were obstacles or de-escalators obstructing contact.
@@
We coded violence when force was used against another’s body (punching, slapping, kicking, hitting, stomping) and/or when another person’s body was forcefully moved. People may defect (not use violence) for all kinds of reasons and due to various causes. To contrast everyone not fighting from conditional cooperators, we coded them as unconditional defectors if others were fighting while they were not. Among them we include de-escalators, as well as group members who were prevented to fight by de-escalators.
We subsumed the following behaviors of members of the opponent group under turmoil for the focal group: aggressing (including fighting gestures); pulling off clothing (jackets or vests); pulling up pants (signaling readiness to fight); pointing toward opponents; provocative gesturing with fingers or hands (as an invitation to engage); bending forward toward an opponent; approaching the focal group; encroaching (invading opponents’ personal space through using or damaging objects belonging to them); teasing, such as lightly hitting or ridiculing; and violence. Because stumbling and falling signal vulnerability of opponents, which tends to agitate focal group members [19, 50, 78], we included these mishaps in our measure. It is likely that over relatively brief time intervals, the effects of turmoil accumulate. Therefore we calculated the level of turmoil from the beginning of the video until a focal group’s (first) maximum participation in violence. We did this by multiplying the duration of each instance of turmoil by the number of individuals involved, and adding up all these weighted instances.
@@
Our data do not enable us to assess causality. Instead, we assess if the patterns in the data support or refute the theoretical predictions.
4.1 Ethics
@@ 5 Results
We confirm the finding of an earlier study [19] that fighting tends to start in small groups or in small subgroups of larger groups. As predicted [10], we found that small (sub)groups burst into action at lower levels of turmoil than larger groups, and lower levels are of course reached earlier. The local emergence (versus central coordination) of collective action in small groups has also been observed in protests [71]. In our data, bursts developed in 13 (37%) of the 35 smallest groups (dyads and triads) and in 10 (42%) of the 24 larger groups. In bursts, the correlation between focal group size and opponents’ turmoil is 0.53. The size-turmoil relationship is slightly disturbed by dyads more often facing a larger opponent group and being therefore less likely to fight collectively than triads, which in turn are more robust in general [66]. [^8] The predicted relation between group size and turmoil is $T\approx n^{1/2}$ [10], but due to the small data and unexplained variance therein, a curve does not fit better than a straight line; hence the reported correlation. If there was a clear winner, it was always the larger group, with only one exception, in line with earlier models of warfare controlling for weapons [38].
@@
The proportions of unconditional defectors in groups with bursts (mean = 0.19; s.d. = 0.21) and groups without bursts (mean = 0.49; s.d. = 0.27) are box-plotted in Figure 4 (Welch test $t=4.796;P=6.411$ x $10^{-6}$; df = 54.76; the ROC-curve has AUC = 0.760). The predicted critical threshold (vertical line in the figure) separates the two boxplots, but one might question how robust this result is, given some noise in coding. There is a risk of missing a violent act (it can be extremely fast) or mislabeling a fast movement as violence, thus mixing up fighters and defectors, and a risk of misassigning individuals to groups. Even if we suppose there is a 10% chance of each of these errors (which corresponds to Krippendorff’s $\alpha=0.75$; see inter-coder reliability in the Supplementary information), the result is hardly weakened (Welch test $t=3.879;P=0.0025$, averaged over 1000 simulation runs). Hence the distinction between bursts and fizzles appears to be robust.
@@
Figure 4: The proportion of unconditional defectors and the predicted critical level (vertical line), which for separates the boxplots of bursts (1) and fizzles (0). The data points are our 59 groups.
@@
The empirical distinction between bursts and fizzles is not perfect, and Figure 4 shows that 15 groups out of 59 (25%) are categorized incorrectly. In 11 of these 15 groups, the time between the first violent act and last participant joining in was above our two-second interval, hence we did not count these collective actions as bursts. In simulations, lags between first instances of violence and the moment when at least half of the group participates frequently occur when turmoil lingers at its critical level (Fig. 5), versus steadily increasing turmoil that yields a clean burst (Fig. 2). Lingering turmoil seems to have occurred in 4 cases, but we would need a better scale of turmoil than the cumulative scale we currently use to be sure. Far more important are local circumstances beyond the model. When re-examining the time lags beyond two seconds, we noticed various causes of delay: (1) group members were in a car and it took time to get out; (2) a focal group member stood at a distance from the fight on a slippery floor next to a pool; (3) a focal group member was beaten and seemed intoxicated by alcohol, hence responded slowly; (4) a focal group member was first pushed over and it took him time to get on his feet again; (5) opponents moved around so fast that it took time to hit one; (6) an opponent fell into thick bushes such that only one focal group member at a time could get to him; (7) group members were sitting in a train and needed time to get on their feet; (8) group members were constrained in a small room where they searched for space to lash out; and, (9) group members were hindered by the mess of two opponents fighting on a collapsed table full of food. One can hardly blame the model for not incorporating these local circumstances that caused delays, and perhaps our two seconds limit should be applied with more lenience in such cases. Finally, two delays were caused by ambiguous behavior: (10) someone tried to de-escalate, was hit, and then attacked, and (11) someone tried to pull back his fighting friend (de-escalation) but simultaneously attacked.
@@
Figure 5: Simulation of Equation 1 at a the critical level of turmoil, $T_{c}\approx 0.11$, over a range of Monte Carlo steps as proxies for time points. The vertical axis shows the proportion of cooperators ($N_{C}/n$) in a fully connected network ($n=5$) with one unconditional defector. Note an extended period of multiple short spikes of violence before all conditional cooperators join in. Similar lags occur without unconditional defectors.
@@
In the 4 out of 15 remaining groups, non-bursts were predicted whereas bursts occurred. Although for parsimony, we applied the same $S_{0}$ (and $\Delta$) for all conditional cooperators, some combatants may actually have a higher value than the average, $S_{0,i}>\overline{S_{0}}$, which could explain these cases.
@@ 5.1 Alternative explanation
Although in 21 out of 23 bursts, some degree of synchronization (10.8 on average) preceded collective violence, there were 18 cases in which synchronization (9.6 on average) was not followed by collective violence. In several of the latter cases, synchrony turned out to be a deceptive performance composed of blustering and aggrandizing without commitment to fighting. This does not imply that synchronization is unimportant (it probably is for solidarity; [19]), but it does not predict collective violence.
6 Discussion and Conclusion
@@
The simple Ising model is a century old [37] and has been applied to a wide range of problems [70, 43], to which we add the dilemma of collective action. It explains cooperation parsimoniously, based on agitating stimuli without recourse to strong rationality, initiative takers, norms, feedback through selective incentives, or reliable information passing through the network, resulting in reputations. The Ising model elucidates the temporal unfolding of violence by predicting a critical threshold of unconditional defectors that distinguishes a burst of collective action from a fizzle, which is largely supported by the data. The model also explains why violent groups are often small or are small subgroups of larger groups despite greater risk. Small (sub)groups have a lower critical threshold of turmoil, and in a confrontation with opponents, lower levels are reached earlier. We also investigated whether synchronous action precedes violence, but we found that synchronization precedes both collective and solitary violence, and cannot predict either of these outcomes. However, synchronization may still be important to increase solidarity [27, 22].
This study has several limitations. Because we did not select videos without violence, we cannot be certain that violence is caused by turmoil. Moreover, developing a proper scale of turmoil or agitation for field studies is exceedingly difficult, due to variations across domains (e.g., street violence versus grievances about politics), and different effect sizes of different stimuli (e.g., killed family members versus a push that might be quickly forgotten). On top of these challenges, our measurements underestimated turmoil because the videos are left-truncated, and our data depend on camera angle and vision width, inevitably excluding instances of turmoil. Measuring turmoil or agitation is only possible in exceptional situations, for example when protest movements produce it by themselves through accelerated posting of online messages. In these particular cases, the inter-posting time intervals follow a certain pattern (Moore’s law), which makes it possible to predict $T_{c}$ ahead of large street demonstrations [33]. In lab experiments, it is unlikely that the patterns of bursts and fizzles would have been discovered, but now they have been pointed out in the field, ethically responsible intergroup conflicts might be examined experimentally, benefiting from controlled circumstances and unambiguous measurement of agitation. Then, causality of turmoil and of the critical level of unconditional defectors might be established or refuted.
Furthermore, there might be error in our coding. If the truth were perfectly known and the coding corrected accordingly, the two boxplots in Figure 4 might shift to a small degree, but, given our robustness check, it is unlikely that this shift would affect our main result.
@@
Another limitation is in the Ising model itself. Despite predicting the threshold of unconditional defectors fairly well, several groups ended up in the non-burst category due to local circumstances (ignored by the model) that caused delays beyond our two second limit. Furthermore, the Ising model does not predict the severity of violence. For future studies, it is important to expand the number and diversity of cases, and to develop better software than is currently available to automatically code videos. Finally, when individuals find themselves more often in similar situations, they will learn, which is easier in smaller groups where they have a larger influence on their payoff [13]. Some will change their strategy, and turn into unconditional defectors [2] who try to exploit other group members and maximize their individual payoff instead of maximizing the group’s payoff. For the model, this would require individuals updating their decision rules at subsequent Monte Carlo steps.
@@
In this first empirical application, we showed that the Ising model can explain the unfolding of violence, and in all likelihood, more discoveries lay ahead. Extensions to norms (as external fields) and noise in actors’ information about others’ behavior (i.e., reputations) have been explored in simulations [9]. In all likelihood, it can be applied to protests [29] and revolts, which break out more often if (rumors say that) a government or its police are weakened [67, 74], analogous to vulnerable individuals in street fights. The model also seems applicable to bystanders collectively helping victims under uncertainty [58], and to lynchings [6]. For example, Nussio [52] argued that lynchings in Mexico can be explained through solidarity combined with peer pressure in a network transmitting reputations meshed with local norms. This argument is perfectly consistent with an expanded Ising model [9], but it lacks the agitation necessary for cooperation to start without centralized organization, as it did. We would argue that the agitation was due to the violation of social norms, of which the lynching victims were accused (e.g., child theft), augmented by the gossip transmitted through the network. In other countries, putative norm violations that agitated crowds and lead to lynchings were rape [12] and blasphemy [3, 60].
@@
Taking a broader perspective on the role of random noise, multiple studies have found that it can solve coordination problems, such as collective responses to opponents’ threats, by shaking a group loose from its suboptimal (e.g., non-cooperative) state [65], but it can also disturb an optimal state. In the Ising model, both can happen, depending on the amount of turmoil: when in a group, turmoil approaches $T_{c}$ from below, random noise in decision making facilitates cooperation, but at high levels, too much noise entails confusion and error.
@@ 7 Data and compilation of data set
One concern about phone-recorded clips is that recorders start filming when the antagonism is already ongoing; our data are thus left-truncated to an unknown degree. We discarded footage that immediately started with physical violence. Another concern is that uploaded videos might be biased toward more spectacular cases. However, our sample shows variety in the forms and severity of violence, ranging from groups who engage in frenzied collective violence to incidents that involve just a few slaps. A third concern is whether the recording influenced the behavior of the assailants. Several videos showed more than one person recording the incident. A host of research indicates that people do not change their behavior substantially in the presence of a camera and mainly focus on what they are doing (reviewed by [35]). This is probably even more the case in antagonistic situations, and because young people are used to being filmed by phones. Additionally, when people do pay attention to the camera, which happened in just one of our clips, it is visible in the recording and can be incorporated in the analysis. (In our case, this concerned behavior coded as turmoil.)